# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v1

@@ Abstract
Text embedding has become a foundational technology in natural language processing (NLP) during the deep learning era, driving advancements across a wide array of downstream tasks. While many natural language understanding challenges can now be modeled using generative paradigms and leverage the robust generative and comprehension capabilities of large language models (LLMs), numerous practical applications—such as semantic matching, clustering, and information retrieval—continue to rely on text embeddings for their efficiency and effectiveness. In this survey, we categorize the interplay between LLMs and text embeddings into three overarching themes: (1) LLM-augmented text embedding, enhancing traditional embedding methods with LLMs; (2) LLMs as text embedders, utilizing their innate capabilities for embedding generation; and (3) Text embedding understanding with LLMs, leveraging LLMs to analyze and interpret embeddings. By organizing these efforts based on interaction patterns rather than specific downstream applications, we offer a novel and systematic overview of contributions from various research and application domains in the era of LLMs. Furthermore, we highlight the unresolved challenges that persisted in the pre-LLM era with pre-trained language models (PLMs) and explore the emerging obstacles brought forth by LLMs. Building on this analysis, we outline prospective directions for the evolution of text embedding, addressing both theoretical and practical opportunities in the rapidly advancing landscape of NLP.
@@
I Introduction
Text embedding learning is a fundamental task in Natural Language Processing (NLP), which aims to extract features from text with the natural languages (including words, sentences, and documents). Formally, given the text $\mathcal{X}$, text embedding learning aims at training a model $f:\mathcal{X}\times\mathrm{\theta}\rightarrow\mathbb{R}^{d}$ on the dataset $D\subset\mathcal{X}$. The continue vector $h=f_{\theta}(x)$ output by $f$ is called as the embedding of the text $x$, which is expected to contain valid information in the input text and perform well on the downstream tasks.
Advanced large language models (LLMs) have recently demonstrated exceptional generalization and transfer capabilities in downstream tasks where generative paradigms are applicable, such as information extraction, text classification, and machine translation. However, not all tasks are suitable for modeling using generative paradigms; tasks such as dense retrieval, semantic text similarity, etc., still require text embedding for computing similarity. In this aspect, LLMs are not different from traditional encoder-only PLMs and do not directly produce high-quality embedding. Therefore, how to obtain high-quality text embeddings remains a key problem.
Fortunately, LLMs have comprehension and generation capabilities far beyond those of previous models, which opens up new opportunities for text representation learning. Specifically, LLMs change the existing landscape in two ways: (1) LLMs can act as data annotators for large-scale, high-quality, and fine-grained text datasets, and (2) LLMs can replace current PLMs as model pedestals for generating higher-quality textual representations. Many works have been devoted to introducing large models for obtaining high-quality text embedding. However, most works focus on a single task, and similar methods have been “proposed” repeatedly in different fields. This leads to a lack of comprehensive and objective understanding of the role played by LLMs in text embedding across various fields. At the same time, the development of LLM as introduced a number of emerging tasks. Examples include long context compression developed to improve the reasoning efficiency of LLM, and embedding inversion tasks to study the information leakage in text. Although the two tasks are developed with different motivations, they share the similar framework and use LLM to understand the information in the text embedding.
This survey focuses on deep representation learning methods, which train the encoders to map texts to a high-dimensional space. We hope this survey will help researchers focusing on different tasks find commonalities in the problem, combining efforts to promote more rapid development in this area. Compared to the existing surveys [1, 2] in this field, this survey covers more comprehensive downstream tasks. They not only include the traditional downstream tasks that require text embeddings (such as semantic text similarity, information retrieval, and text clustering) but also contain the emerged tasks in the era of LLMs (such as long context compression and embedding inversion). In addition, this survey discusses the relationship between LLMs and text embedding for the first time, mainly including (1) LLM-augmented text embedding, (2) LLM as text embedders, (3) Text embedding understanding with LLMs. Fig. 1 shows the relationship between the text embedding model (which we call text embedder) and the LLM in each section.
The subsequent sections will unfold in the following order: Section II introduces the developed stage, the training and evaluation tasks for text embedding; Section III introduces the text embedding methods where LLMs are used for data augmentation and while another model is used to be the text embedder; Section IV presents the text embedding methods where LLMs are the backbones of text embedders; Section V show two novel text embedding-related tasks in the LLM era: long context compression and embedding inversion. Section VI discusses the remained challenge before the presentation of LLMs and the emerging challenge in the era of LLMs; Section VII shares several promising trends and direction in the text embedding field recently.
II Priliminary
II-A History of Text Embedding
II-A1 Era of Statistics Machine Learning
@@
In the early days, researchers primarily relied on manually designed features to represent text. The development of the method is accompanied by a transformation of the output vector’s form, which is from bag-of-words models (one-hot vectors) to TF-IDF [3] (sparse vectors), then to word embeddings (dense vectors). These methods require domain experts to select and design features manually, and the quality and quantity of features limit their effectiveness. As machine learning techniques advanced, researchers began exploring methods to learn text embeddings using statistical approaches such as Latent Semantic Analysis (LSA) [4], Latent Dirichlet Allocation (LDA) [5], and others. While these methods can automatically learn text embedding, they still have limitations in capturing complex semantic and syntactic structures.
II-A2 Era of Deep learning
@@
With the emergence of deep learning techniques, word embedding has become a significant breakthrough in text embedding learning. Models such as Word2Vec [6], GloVe [7], and FastText [8] can map words into a low-dimensional continuous vector space to capture semantic and syntactic relationships between words. Over the past few years, pre-trained language models (PLMs) with millions of parameters (e.g., BERT [9] and RoBERTa [10]) have been proposed first. The “pretraining then fine-tuning” paradigms have performed well on various downstream tasks and played an important role in practical on-the-ground applications. However, the embedding spaces of these PLMs are proved to be anisotropic [11], leading to surprisingly high similarities computed for any two texts.
In this phase, the centre of text embedding learning shifts to fine-tuning PLMs for higher-quality text embeddings. Some supervised methods were explored first, followed by post-processing and unsupervised methods. Despite the tremendous progress, there is still a huge performance gap between supervised and unsupervised methods. Then contrastive learning is introduced to use the negative example fully, which focuses on constructing positive and negative examples for the samples in each dataset and training the model to differentiate between positive and negative examples. The common contrastive learning loss is InfoNCE [12], which can be expressed as
where $s:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the distance function, while $x^{+}$ and $\{x^{-}_{j}\}_{j=1}^{N}$ is the positive example and negative examples for the text $x$, respectively. Improvements in contrastive learning have focused on (1) the form of the loss function [13, 14, 15, 16] and (2) better methods for constructing positive and negative samples [17, 18, 19, 20].
II-B Large Language Model
We use the term “large language models” or “LLMs” to exclude pre-trained language models based on the Transformer encoder (such as BERT and RoBERTa) and neural networks with typically smaller parameter counts (such as LSTM [21] and GRU [22]). We refer to the language models whose variants’ parameter number is more than 1B as LLMs. Therefore, the LLMs we study include:
TABLE I: Representative LLMs contributing to advancements in text embedding.
Type Representives
Encoder-Decoder T5 [23], FLAN-T5 [24]
Decoder-Only GPT-Neo [25] , BLOOM [26], OPT [27]
Mistral [28], LLaMA Series [29, 30, 31]
Vicuna [32], Qwen Series [33, 34, 35]
Commercial API OpenAI GPT [36], Gemini [37]
II-C Training Dataset
@@
Text-embedded data involves multiple downstream applications, and advanced methods often integrate massive amounts of data from various domains in various languages involved in training. Table II shows non-synthetic datasets that are often used in existing work. Given the enormous amount of work and the large differences between methods, we will not present the training data used by each method. In addition, current works tends to use zero-shot setting to test the generalization ability of text embeddings, so the practice of comparing methods with different training data on the same benchmark has been widely accepted by the community.
II-D Downstream Task for Evaluation
Text embedding learning often serves multiple downstream tasks and cannot be trained using label $Y$ as in text classification tasks. Therefore, one research focus of text embedding learning mainly lies in designing pretext tasks and constructing corresponding datasets.
II-D1 Semantic Textual Similarity
The Semantic Textual Similarity (STS) task quantifies the semantic similarity degree between two texts.
Evaluation Dataset
The dataset for evaluating STS can be expressed as $D_{\rm eval}=\{(x_{i}^{a},x_{i}^{b},c_{i})\}_{i=1}^{n}$, where $(x_{i}^{a},x_{i}^{b})$ is a text pair and $c_{i}$ is the average similarity between the two texts. The popular evaluation protocol usually contains seven STS datasets for English setting, which contains STS12-16 [38, 39, 40, 41, 42], STS-B [43] and SICK-R [44] in SentEval Benchmark [45] and STS-17 [43] and STS-22 [46] for multi-lingual setting.
Evaluation Metric
The model is expected to take two texts $(x_{i}^{a},x_{i}^{b})$ as input and produce a predicted score $\hat{c}_{i}$ that represents the similarity of the two texts. Common evaluation metrics include the Pearson correlation coefficient and the Spearman correlation coefficient. The Pearson correlation coefficient primarily assesses the correlation between absolute values of $c_{i}$ and $c_{i}^{p}$, which is expressed as
where $\bar{c}_{i}$ and $\bar{c}_{i}^{p}$ are the mean value of $\{c_{i}\}_{i=1}^{n}$ and $\{c_{i}^{p}\}_{i=1}^{n}$, respectively. And the Spearman correlation coefficient primarily assesses the correlation between the ranks of $c_{i}$ and $c_{i}^{p}$ in their respective lists, which is expressed as
$r_{\rm spearman}=1-\frac{6\sum_{i=1}^{n}(r_{i}-r_{i}^{p})^{2}}{n(n^{2}-1)}$ (3)
where $r_{i}$ and $r_{i}^{p}$ are the rank of $c_{i}$ and $c_{i}^{p}$ in their respective lists, respectively.
II-D2 Information Retrieval
Information retrieval (IR) aims to retrieve documents relevant to a query, where their matching scores measure the relevance between the query and the documents. (In this survey, we focus on text retrieval.) Information retrieval systems can be roughly divided into two stages. The first stage (retriever) aims to reduce the candidate space by retrieving relevant candidate documents. The second stage (reranker) mainly focuses on reranking the candidate documents.
Evaluation Dataset
Generally, the dataset for training and evaluation can be written as $D=\{q_{i},d^{+}\}_{i=1}^{n}$, where $q_{i}$ is the query and $d^{+}$ are the positive documents that match $q_{i}$. The popular evaluation protocol usually use BEIR [47] for English setting, MIRACL [48] and Mr.TyDi [49] for multi-lingual setting.
@@ II-D2 Information Retrieval
TABLE II: Detailed information of available fine-tune datasets. Here, we categorize the available fine-tuned datasets into five major tasks, namely information retrieval(IR), question answering(QA), natural language inference(NLI), classification and multi-task (Multi).
@@
AmazonCounter [69] Multi E-commerce Classification 24,000 - 24,000
Emotion [70] English Twitter Classification 16,000 - 16,000
MTOPIntent [71] Multi Web Classification 18,800 - 18,800
ToxicConversationsClassification  [^2] https://www.kaggle.com/competitions/jigsaw-unintended-bias-in-toxicity-classification/overview English Twitter Classification 50,000 - 50,000
TweetSentimentExtraction  [^3] https://www.kaggle.com/competitions/tweet-sentiment-extraction/overview English Web Classification 27,500 - 27,500

### v2

@@ Abstract
Text embedding has become a foundational technology in natural language processing (NLP) during the deep learning era, driving advancements across a wide array of downstream tasks. While many natural language understanding challenges can now be modeled using generative paradigms and leverage the robust generative and comprehension capabilities of large language models (LLMs), numerous practical applications—such as semantic matching, clustering, and information retrieval—continue to rely on text embeddings for their efficiency and effectiveness. Therefore, how to combine the LLMs and the text embeddings has become one of the hotspots of academic attention in recent years. In this survey, we categorize the interplay between LLMs and text embeddings into three overarching themes: (1) LLM-augmented text embedding, enhancing traditional embedding methods with LLMs; (2) LLMs as text embedders, adapting their innate capabilities for high-quality embedding; and (3) Text embedding understanding with LLMs, leveraging LLMs to analyze and interpret embeddings. By organizing recent works based on interaction patterns rather than specific downstream applications, we offer a novel and systematic overview of contributions from various research and application domains in the era of LLMs. Furthermore, we highlight the unresolved challenges that persisted in the pre-LLM era with pre-trained language models (PLMs) and explore the emerging obstacles brought forth by LLMs. Building on this analysis, we outline prospective directions for the evolution of text embedding, addressing both theoretical and practical opportunities in the rapidly advancing landscape of NLP.
1 Introduction
Text embedding learning is a fundamental task in Natural Language Processing (NLP), which aims to extract features from text with the natural languages (including words, sentences, and documents). Formally, given the text $\mathcal{X}$, text embedding learning aims at training a model $f:\mathcal{X}\times\mathrm{\theta}\rightarrow\mathbb{R}^{d}$ on the dataset $D\subset\mathcal{X}$. The dense vector $h=f_{\theta}(x)$ output by $f$ is called the embedding of the text $x$, which is expected to contain valid information in the input text and perform well on the downstream tasks.
Advanced large language models (LLMs) have recently demonstrated exceptional generalization capabilities in downstream tasks where generative paradigms are applicable, such as information extraction, text classification, and machine translation. However, not all NLP tasks are suitable for modeling with generative paradigms; tasks such as dense retrieval, semantic text similarity, etc., still require text embedding for computing similarity. The powerful semantic understanding capability of the LLMs reveals the ability to use it to obtain high-quality embeddings. However, decoder-only LLMs are not different from traditional encoder-only pre-trained language models (PLMs), facing the anisotropy challenge in the native embedding space [1]. It is manifested by the high similarity of two token embeddings, which does not reflect their semantic similarity. Higher-level text, such as sentences and documents, suffer from similar problems since they often share embedding space with tokens [2]. Fortunately, LLMs have comprehension and generation capabilities far beyond previous PLMs, which opens up new opportunities for text-embedding learning. Specifically, LLMs change the existing landscape in two ways: (1) LLMs can act as data annotators or generators for large-scale, high-quality, and fine-grained text datasets, and (2) LLMs can replace current PLMs as the backbone for generating higher-quality text embeddings. Many works have been devoted to introducing LLMs to obtain high-quality text embedding; however, they focus on a single downstream task, and similar methods have been proposed repeatedly in different research fields. This leads to a lack of comprehensive and objective understanding of the role played by LLMs in text embedding research.
At the same time, the development of LLM has introduced many emerging tasks, some of which are highly relevant to text embedding. This survey introduces two novel downstream tasks with the active community: long context compression and text embedding inversion. Long text compression (ICC) aims to compress long text into a compact number of embeddings while preserving key information. Compared to the context length extension, ICC can improve the decoding efficiency of LLMs and has great potential for application in paradigms such as retrieval-augmented generation (RAG). Embedding inversion aims to recover the original text information from its embedding. With the rapid development of vector databases and commercial embedding services, the study of embedding inversion is significant for protecting privacy and security in embedding. Although the two tasks are developed with different motivations, they share a similar learning framework and use LLM to understand the information in the text embedding.
This survey focuses on deep embedding-learning methods in the LLM era, which include the latest works focused on the traditional downstream tasks related to text embeddings (such as semantic text similarity, information retrieval, and text clustering) and the pioneering work focused on novel downstream tasks (such as long context compression and embedding inversion). Specifically, this survey summarizes the three relationships between LLMs and text embedders for the first time, mainly including (1) LLM-augmented text embedding, (2) LLM as text embedders, and (3) Text embedding understanding with LLMs (Fig. 1), while the existing related surveys [3, 4, 5] usually covers only limited works on LLMs embedders. We hope this survey will help researchers in different communities find commonalities in the problem, combining efforts to promote more rapid development of text embedding with the help of LLMs.
The subsequent sections will unfold in the following: Section 2 introduces the developed stage, the training and evaluation tasks for text embedding; Section 3 introduces the text embedding methods where LLMs are used for data augmentation and while another model is used to be the text embedder; Section 4 presents the text embedding methods where LLMs are the backbones of text embedders; Section 5 show two novel text embedding-related tasks in the LLM era: long context compression and embedding inversion. Section 6 discusses the remaining challenge before the presentation of LLMs and the emerging challenge in the era of LLMs; Section 7 shares several promising trends and directions in the text embedding field recently.
@@ 1 Introduction
2 Priliminary
2.1 Brief History of Text Embedding
2.1.1 Era of Statistics Machine Learning
@@
Text embedding is first applied in information retrieval, text mining, and machine translation. In the early days, researchers primarily relied on manually designed features to represent text, which are used to measure the relation between the texts. These methods require domain experts to select and design features carefully, and the quality and quantity of features limit their effectiveness. The development of the method is accompanied by a transformation of the output vector’s form, which is from bag-of-words models (one-hot vectors) to TF-IDF [6] (sparse vectors), then to text embeddings (dense vectors). As machine learning techniques advanced, researchers began exploring methods to learn text embeddings using statistical approaches such as Latent Semantic Analysis (LSA) [7] for information retrieval and Latent Dirichlet Allocation (LDA) [8] for topic mining. Though these methods can automatically learn low-dimensional text embedding, there are still obvious shortcomings. For example, LSA is hard to capture complex semantic and syntactic structures, while LDA is difficult to apply to large-scale datasets.
2.1.2 Era of Shallow Neural Networks
With the development of neural machine learning, the NLP community discovered that well-learned embeddings can be transferred to many downstream tasks [9, 10, 11]. Therefore, word embedding (distributed representation [12]) is widely studied as a fundamental technique in NLP instead of a subsidiary to other tasks. We introduce representative text embedding methods in this era from three perspectives: training data, model architecture, and learning paradigm.
Training Data
The training data for word embedding mainly includes two kinds: (1) large-scale text corpus for unsupervised training and (2) high-quality semantic data for secondary supervised training. For example, Wikipedia dump  [^1] https://dumps.wikimedia.org/, Gigaword5  [^2] https://catalog.ldc.upenn.edu/LDC2011T07 and Common Crawl  [^3] https://commoncrawl.org/ are usually used for large-scale training corpus. Based on learned word embedding, WordNet [13] and Paraphrase Database (PPDB) [14] are used for semantic knowledge enhancement.
Model Architecture
Though deep neural networks [15, 16] have shown amazing potential in the same time period, considering the massive size of the corpus, shallow neural network, such as a single-layer Feedforward Neural Network (FNN), is used to maps each word to the embedding.
Learning Paradigm
With the emergence of deep learning techniques, word embedding become a significant breakthrough in the NLP community. (1) Unsupervised paradigm. For word-level, Word2Vec [17] proposes Continuous Bag of Words (CBOW) for predicting the center word based on the context and Skip-Gram for predicting the context based on the center word. Word2Vec has two drawbacks: (a) inability to model global statistical Information; (b) inability to handle out-of-vocabulary (OOV) words and the internal structure of words, and these drawbacks are improved by GloVe [18] and FastText [19], respectively. Specifically, GloVe utilizes the global information adequately by pre-constructing the global word co-occurrence matrix, showing that high-quality word embeddings can be obtained without neural networks. FastText represents each word as a collection of n-grams of characters, thus capturing the word’s internal structure, which helps to model OOV words. Beyond word-level embedding, Doc2Vec [20] proposes Distributed Memory (DM) and Distributed Bag of Words (DBOW) to extend embedding to the document level (e.g., sentences, paragraphs, or entire documents). SIF [21] shows that the sentence embedding obtained by weighted averaging of the word embeddings can be regarded as a strong baseline. Besides, removing the projection on the first principal component for each embedding [21, 22] will increase the differentiation among these embeddings. (2) Supervised paradigm. Syntactic and semantic knowledge are widely introduced as the supervised signals to improve Word2Vec. For example, DEPS embedding [23] replaces the sliding-window style contexts with dependency-based, where only the modifiers can be the context for the target word. RCM [24] and Paragram embedding [25] introduce the semantic relation in Paraphrase Database and WordNet as the constraint to learning the word embedding can reflect these relations.
2.1.3 Era of Deep Neural Networks
With the development of deep learning, embedding methods with deep neural networks (DNNs) become mainstream and demonstrate remarkable performance on a variety of downstream tasks. Compared to word-level embedding, sentence-level embeddings will be more useful in downstream tasks. Although the simple weighting of word vectors is a strong baseline, it is still worthwhile to explore further how to obtain high-quality sentence embedding on learned word embedding.
Training Data
The labelled datasets for NLP fundamentals and downstream tasks have gradually scaled up, providing the opportunity to learn high-quality text embedding. At the same time, DNNs can be trained on the unlabeled corpus with billions of words with the development of GPUs. Therefore, various datasets for basic NLP research and downstream datasets are widely explored for learning to embed. For example, BookCorpus [26] and the parallel corpus from WMT14 [27] and WMT15 [28]; in addition, some popular datasets widely used for embedding learning, such as MS-MARCO [29], SNLI [30], and MNLI [31] are proposed.
Model Architecture
Two architectures that can be summarized, which mainly depend on the training task. (1) Encoder-Decoder for generative training tasks. The generative tasks, such as machine translation [32] and adjacent sentence generation [33], are beneficial for generalizing text embeddings. The backbones of decoder can be RNN [34], LSTM [33], GRU [35] and Transformer [36], while the backbones of encoder can be CNN [37] and DAN [36] besides all above models; (2) Dual-Encoder for discriminative training tasks. The model trained on some discriminative tasks, such as paraphrase matching [38], natural language inference (NLI) [39], and adjacent sentence prediction [40], shows the potential of universal sentence embedding in multiple downstream tasks. Except for the standard DNNs [38, 39], many customized architectures [41, 42] and pooling strategy [43] have been explored for training text embeddings.
Learning Paradigm
Three main lines of work for the training paradigm can be summarized: (1) Unsupervised paradigm. The works for unsupervised embedding learning focus on the design of pretext tasks on the unlabeled data. Skip-Thoughts [33] uses an encoder-decoder architecture that encodes the target sentence and then decodes it to generate the adjacent sentences to learn the embedding. FastSent [44] improves the efficiency of Skip-Thoughts with two efforts: (a) using the sum of word embedding as the sentence embedding for faster encoding speed, and (b) ignoring word order of the adjacent sentences when prediction for fast training speed. Quick-Thoughts [40] further proposes a discriminative task, training the embedding to select the sentence adjacent to the current sentence in a candidate sentence set. (2) Supervised paradigm. The works under the supervised paradigm focus on finding the specific training task and data on which the embeddings acquired after training can be effectively generalized to other tasks. For example, InferSent [39] finds that the embeddings learned on natural language inference (NLI) data have excellent transfer performance in other downstream tasks, while CoVe [32] finds that the word embedding learned by neural machine translation (NMT) can generalize well to other tasks. (3) Multi-task paradigm. Some works, such as GenSen [35] and USE [36], make use of the successful practical experience in each setting, both of which introduce multiple tasks in the previous two paradigms and obtain the sentence embedding for universal tasks.
2.1.4 Era of Pre-trained Language Models
@@
Over the past few years, pre-trained language models (PLMs) with millions of parameters have been proposed first. The “pretraining then fine-tuning” paradigms have performed well on various downstream tasks and played an important role in practical on-the-ground applications. However, the word embeddings of the vanilla Transformer [45] and these PLMs [1] are proved to be concentrated in high-dimensional conical space, leading to surprisingly high similarities computed for any two words. Therefore, the research centre of text embedding learning shifts to improve the embedding space of PLMs.
Training Data
@@
Text-embedded data involves multiple downstream applications, and advanced methods often integrate massive amounts of data from various domains in various languages involved in training. Table 1 shows non-synthetic datasets often used in existing work. Given the enormous amount of work and the large differences between methods, we will not present the training data used by each method. In addition, current works tend to use zero-shot or few-shot settings to test the generalization ability of text embeddings, so the community has widely accepted the practice of comparing methods with different training data on the same benchmark.
@@
Table 1: Detailed information of available training datasets. The datasets are divided into five major categories, namely information retrieval (IR), question answering (QA), natural language inference (NLI), classification and multi-task (Multi).
Dataset Language Domain Task Text-Text / Text-Label Pair
SNLI [30] English Web NLI 550,152
MNLI [31] English Web NLI 392,702
AmazonCounter [46] Multi E-commerce Classification 24,000
Emotion [47] English Twitter Classification 16,000
MTOPIntent [48] Multi Web Classification 18,800
ToxicConversationsClassification  [^4] https://www.kaggle.com/competitions/jigsaw-unintended-bias-in-toxicity-classification/overview English Twitter Classification 50,000
TweetSentimentExtraction  [^5] https://www.kaggle.com/competitions/tweet-sentiment-extraction/overview English Web Classification 27,500
@@
Model Architecture
Most embedding learning works in the era are based on BERT [68] and RoBERTa [69], which consist of multi-layer Transformer Encoders and pre-training on tens of billions of words ($\sim$3.3B words for BERT and $\sim$19B words for RoBERTa).
Learning Paradigm
(1) Post-hoc paradigm. Pioneering attempts to focus on post-processing methods to improve the quality of the embeddings. For example, adjusting the embedding space to an isotropic space by learning the flow function [2] or transforming with whitening matrix [70] has proved effective. Besides, standardizing a few undesirable dimensions of the embeddings [71] is also an effective post-processing method. (2) Unsupervised & Supervised fine-tuning paradigm. After a brief period of exploration for other loss function [72], the supervised and unsupervised paradigms gradually unify to contrastive learning [73, 74, 75, 76, 77]. For each sample (“anchor”) in the dataset, Contrastive learning [78] focuses on constructing positive and negative examples and optimizing the embedding space to close the anchor-positive distance while large the anchor-negative distance. The difference between the supervised and unsupervised settings is mainly reflected in the construction of positives. In the supervised setting, the anchor-positive pair can be constructed based on the existing labeled dataset, such as query-document pair in the retrieval dataset [74] and hypothesis-entailment pair in the NLI dataset [73], etc. In the unsupervised setting, (a) two data-augmented views of the same text [76] or (b) two adjacent texts in the same document [75] are regarded as an anchor-positive pair. The augmentation can be literal-level (such as word delete [73] and back translation [79], etc) or embedding-level (such as dropout [76]). Negatives can be obtained by random sampling in the dataset, and the hard negative mining methods [80] can be used to discover those challenge negatives with more help for embedding learning. The typical loss function of contrastive learning is InfoNCE [78]. Given the anchor sample $x$, the positive example $x^{+}$ and the negative examples $\{x_{j}^{-}\}_{j=1}^{N}$, denote their $d$-dimensional embedding as $h$, $h^{+}$ and $\{h_{j}^{-}\}_{j=1}^{N}$, separately, then the InfoNCE Loss is expressed as
where $s:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the distance function, while $x^{+}$ and $\{x^{-}_{j}\}_{j=1}^{N}$ is the positive example and negative examples for the text $x$, respectively. Improvements based on the contrastive learning have focused on (a) the form of the loss function [81, 82, 83, 84] and (b) better methods for constructing positive and negative samples [79, 85, 86, 80]; (3) Pre-training paradigm. Contrastive learning is useful on large-scale corpora but takes big performance hits on low data situations [74]. This observation shows that the pre-training tasks of PLMs, i.e., Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), etc., do not allow the PLMs to adapt to semantic aggregation quickly and output high-quality text embeddings. Therefore, some work has been devoted to designing better pre-training tasks to improve the performance of contrastive learning in low-data scenarios. Early explorations [87, 88] used only additional incremental pre-training tasks, such as the Inverse Cloze Task (ICT), to improve contrast learning performance. Subsequent improvements involve the modification of model architectures. For example, Condenser [89] restores the masked token using the [CLS]’s hidden states from the last layer and other tokens’ hidden states from a shallow layer. Therefore, the newly generated information in the deep layers has to aggregate in the [CLS] to reconstruct the masked information. SEED-Encoder [90] and RetroMAE [91] feed the final hidden state of [CLS] into an auxiliary decoder to restore the masked information, encouraging the aggregation of information at [CLS].
2.2 Large Language Model
We use the term “large language models” or “LLMs” to exclude encoder-only PLMs (such as BERT and RoBERTa) and deep neural networks with typically smaller parameter numbers (such as LSTM [92] and GRU [93]). We refer to the language models whose parameter number is more than 1B as LLMs, which contain large-size encoder-decoder-based PLMs and decoder-only PLMs. We also take into account commercial APIs that are explicitly LLM-backed behind them. Therefore, the LLMs we study can be summarized in Table 2.
Table 2: Representative LLMs contributing to advancements in text embedding.
Type Representives
Encoder-Decoder T5 [94], FLAN-T5 [95]
Decoder-Only GPT-Neo [96] , BLOOM [97], OPT [98]
Mistral [99], LLaMA Series [100, 101, 102]
Vicuna [103], Qwen Series [104, 105, 106]
Commercial API OpenAI GPT [107], Gemini [108]
2.3 Evaluation Task
The general trend in text embedding is toward cross-task generalization, and thus, mainstream evaluation protocols tend to evaluate an increasing variety of downstream tasks. Here, we introduce the classical downstream tasks used to evaluate embedding quality. Specifically, we give the definition, common evaluation datasets and evaluation metrics for five tasks, including Semantic Textual Similarity (STS), Information Retrieval (IR), Universal Embedding (UE), Long Context Compression (LCC) and Text Inversion (TI).
2.3.1 Semantic Textual Similarity
Semantic Textual Similarity (STS) task [109] is the task to judge the similarity calculated by embeddings whether matching the ratings from humans. Strictly speaking, STS is not a downstream task with a well-defined application scenario, but rather an intrinsic evaluation task that was purposely presented [110]. The motivation behind it is that the distance between two texts’ embedding should reflect the semantics similarity between these two texts.
Evaluation Dataset
The dataset for evaluating STS can be expressed as $D_{\rm eval}^{\rm STS}=\{(x_{i}^{a},x_{i}^{b},c_{i})\}_{i=1}^{n}$, where $(x_{i}^{a},x_{i}^{b})$ is a text pair and $c_{i}$ is the average similarity between the two texts. The popular evaluation datasets usually contains seven STS datasets for English setting, which contains STS12-16 [109, 111, 112, 113, 114], STS-B [110] and SICK-R [115] in SentEval Benchmark [116] and STS-17 [110] and STS-22 [117] for multi-lingual setting.